### v1

# Privacy Technologies for Financial Intelligence

Abstract
Financial crimes like terrorism financing and money laundering can have real impacts on society, including the abuse and mismanagement of public funds, increase in societal problems such as drug trafficking and illicit gambling with attendant economic costs, and loss of innocent lives in the case of terrorism activities. Complex financial crimes can be hard to detect primarily because data related to different pieces of the overall puzzle is usually distributed across a network of financial institutions, regulators, and law-enforcement agencies and they cannot be easily shared due to privacy constraints. Recent advances in Privacy-Preserving Data Matching and Machine Learning provide an opportunity for regulators and the financial industry to come together to solve the risk-discovery problem with technology. This paper provides a survey of the financial intelligence landscape and where opportunities lie for privacy technologies to improve the state-of-the-art in financial-crime detection.
1 Introduction
Our economy depends on a well functioning banking and financial system that has the trust of the general population. However in recent years, the reputation of some major banks has been severely damaged as criminal elements have exploited systemic weaknesses to launder huge sums of money, including transfers to terrorist organisations [Lyn22, LF23]. Besides reputational damage to financial institutions, complex financial crimes like terrorism financing and money laundering can have real direct impacts on society too, including
  • abuse and mismanagement of public funds;
  • increase in societal problems like drug trafficking, illicit trade and gambling, people smuggling, with attendant economic costs; and
  • loss of innocent lives in the case of terrorism activities.
Complex financial crimes can be hard to detect primarily because data related to different pieces of the overall puzzle are usually distributed across a network of financial institutions, regulators, and law-enforcement agencies [LCSH19]. The problem is also rapidly increasing in complexity because of new emerging value-transfer platforms like crypto-currencies and real-time settlement systems.
In early 2016, AUSTRAC, Australia’s Financial Intelligence Agency, set up the Fintel Alliance [^1] https://www.austrac.gov.au/partners/fintel-alliance, an organisation that includes AUSTRAC and the major financial institutions and law-enforcement agencies, to improve public-private collaboration in combating financial crimes. Since then, financial intelligence agencies around the world have started building public-private data-sharing platforms to share and disrupt criminal entities engaging in financial crimes, including the Joint Money Laundering Intelligence Task Force (JMLIT) [^2] https://www.nationalcrimeagency.gov.uk/what-we-do/national-economic-crime-centre in UK, and the Financial Crime Enforcement Network Exchange [^3] https://www.fincen.gov/resources/financial-crime-enforcement-network-exchange in the United States. These public-private platforms are highly effective against known threats.
However, their discovery functions to find as-yet unknown risks in the financial system are still not as strong as they can be, primarily because of privacy constraints around large-scale data matching and analytics between organisations. Recent advances in Privacy-Preserving Data Matching and Machine Learning provide an opportunity for regulators and the financial industry to come together to solve the risk-discovery problem with technology. This paper provides a survey of the financial intelligence landscape and where opportunities lie for privacy technologies to improve the state-of-the-art in financial-crime detection. Our paper complements the material covered in [Max20], which explored global case studies of public and private sectors employing cryptographic-based privacy technologies to foster collaborative financial intelligence. [^4] https://www.future-fis.com/the-pet-project.html
2 Financial Intelligence
Financial intelligence is a highly specialised field that is not well understood by many. We provide an introduction to the financial intelligence ecosystem in this section, including a description of the current state, some existing limitations and opportunities.
2.1 The Financial Intelligence Ecosystem
Financial Intelligence Units (FIUs) around the world are governed by anti-money laundering and counter-terrorism financing (AML/CTF) regimes that are quite similar in nature [Mad11, Gil04]. Generally speaking, FIUs collect data like threshold transaction reports, international fund transfer instructions, and suspicious matter/activity reports from Reporting Entities (REs), which include banks, money remitters, casinos, law firms, real-estate companies, and financial companies. (In a country the size of Australia, there are about 10,000 REs in the financial system.) FIUs may also get data about entities of interest from partner agencies like law-enforcement agencies and partner FIUs in other countries. Using all these collected data, analysts in FIUs produce intelligence products that either report on specific threats that can be actioned by law enforcement, or describe macro threats and mitigation strategies that financial institutions should adopt in their internal controls. Figure 1 shows the general information flow from reporting entities to FIUs and, from there, to law-enforcement agencies.
Figure 1: A reporting entity is a company that provides financial services that are regulated under AML/CTF laws; KYC = Know Your Customer; ECDD = Enhanced Customer Due Diligence; TTRs = Threshold Transaction Reports; IFTIs = International Fund Transfer Instructions; SMRs = Suspicious Matter Reports; API = Application Programming Interface.
There are safeguards and risk monitoring all over the ecosystem, inside each reporting entity and within the FIU, and there is a reasonable mix of automation and human involvement in those risk-monitoring and risk-reduction processes. The parts of the diagram that are painted in brown are areas of the financial intelligence ecosystem where advanced analytics like predictive and prescriptive modelling can play an important role, mostly because the data volume that need to be processed at those places are enormous and the relationships that need to be mastered and managed, including cause and impact relationships, are often complex.
Generally speaking, all the data received by an FIU are pushed through an entity resolution step to identify and merge records that are highly likely to refer to the same entity [ZNCC18]. This is usually a highly non-trivial computational problem because of the sheer number of parties involved in financial transactions (billions for a typical FIU), to arrive at a table like Figure 2, where the rows are entities (either individuals, companies, or groups of related individuals/companies) and the table show their behaviour in the financial system as seen through the lens of different reporting entities. In some cases, we also have explicit risk indicators for these entities. Given such a table, an FIU equipped with data science expertise and big data analytics platforms can use profiling and statistical machine learning techniques, both supervised and unsupervised at different levels of granularity from individual transactions to single entities and networks of entities, to uncover risks in the financial system and act on them. Examples of such work include [SWC+16, SWZ+17, SZC+17].
Figure 2: The entity-resolved data and risk indicators available to a typical FIU.
A key thing to understand about financial intelligence is that financial crimes cannot really be detected from financial transaction data alone. Here, in a nutshell, are the sort of things we can infer about anomalies in financial transactions [Mad11, RT04]:
  1. 1.
    the source of fund cannot be explained;
  2. 2.
    there are rapid movements of funds through different channels, which may indicate intention to hide the money trail;
  3. 3.
    the financial transaction behaviour is inconsistent with a customer’s expected business profile.
These sorts of activities are not illegal on their own, of course. It is only when these activities accompany a predicate crime — tax evasion, welfare fraud, drug trafficking, etc — that criminality can be established and treated. For that reason, one needs law-enforcement presence throughout the ecosystem to supply leads and the risk contexts for the whole system to work well. In particular, pure discovery work in financial transaction data is quite difficult but commercial entities of interest databases like LSEG World-Check and external datasets like company registration data can be useful in adding context to financial data.
That is the current state. The overall system is far from perfect but, by and large, the regime works reasonably well in deterring, detecting, and disrupting financial crimes. The main limitation is in the links that are not there in the ecosystem. First of all, a lot of financial transaction data, including domestic interbank transfers regardless of amount, are not reported by default to an FIU. Secondly, there are literally tens of thousands of reporting entities in a mature economy like Australia and these reporting entities are largely limited in what intelligence and data they can share with each other because of legal impediments around sharing SMRs between reporting entities to minimise the risk of (accidental) tipping off. There is also not a lot of feedback loops built into the system, including that between reporting entities and an FIU, and that between an FIU and law-enforcement agencies.
The important limitations in the current state can thus be mostly traced to the following two issues:
  • Complex financial crimes typically involve transactions that involve multiple financial institutions in multiple geographies, going through multiple payment channels, some of which are unregulated.
  • FIUs and reporting entities are all working with data silos and no one has a complete view of activities across the financial system.
To appreciate the extent of the problem, one only needs to look at the difficulty of trying to answer a simple question like finding all entities that exhibit a certain cross-border transaction pattern like that shown in Figure 3.
Figure 3: A simple hard problem
The question, and many more complicated others we would like to ask, simply cannot be answered today because the data needed to answer them sit in different organisations in different countries, governed by different legislation.
2.2 Privacy Technologies As A Way Forward
So what can we do? Figure 4 shows a possible way forward using distributed confidential computing technologies, assuming FIUs and partner agencies are willing to work together but without needing to make wholesale changes to AML/CTF legislations everywhere.
Figure 4: Privacy-enhanced version of Figure 2. LEAs = Law Enforcement Agencies; PAs = Partner Agencies; EOI = Entities of Interest
The obvious obstacle to different FIUs and partner agencies sharing large volumes of data with each other is privacy concerns. In the proposed scheme, instead of entity resolution, we have all the participating agencies contributing data that are first matched using privacy-preserving record linkage (PPRL) algorithms [CRS20]. In a PPRL algorithm, the records in different databases that represent or belong to the same entity are identified in a privacy preserving manner without revealing any plaintext values to the participating parties. The matched data are then encrypted using a homomorphic encryption scheme [Hal17] (as described in Section 3.2), which allow mathematical operations to be performed on encrypted data, to arrive at a table similar to what we showed earlier, with two distinctions:
  • The values in the table are all encrypted and the columns of the table are stored in a distributed way across different databases in different agencies.
  • The point of using homomorphic encryption schemes is that it is still possible to run a range of risk-modelling algorithms over the encrypted data, and once we get to a small set of entities of interest, decrypting those entities’ data can be done without raising any privacy concerns for the general public.
Having the data distributed in different data sources does complicate the running of risk-modelling algorithms somewhat but the added security and scalability make it a worthwhile trade-off to make.
There are basically three fundamental computational problems to solve, in increasing order of difficulty:
  1. 1.
    How can two FIUs, each of which has a list of suspicious entities, share only the common suspicious entities with each other, without sharing their entire list with each other? More generally, how can such sharing be done between multiple FIUs?
  2. 2.
    Given a criminal typology that can be described as a graph pattern, where nodes denote entities and edges denote relationships, how can we detect all entities that satisfy the criminal typology when the underlying graph span multiple financial institutions?
  3. 3.
    How can we perform model learning, either supervised or unsupervised, over a set of entities whose financial data and behaviour are spread across multiple financial institutions?
In the rest of this paper, we will look at how aspects of that future state can be realised through different privacy technologies.
3 Privacy Technologies
The purpose of this section is to briefly introduce the most common privacy technologies that are applicable to financial intelligence. These techniques are based on different mathematical foundations to serve different data privacy and security purposes. Secure multiparty computation, through the use of secret sharing schemes, permits multiple parties to jointly compute a function without divulging each party’s secret information to the others. Homomorphic encryption tackles this issue from a different perspective, by encrypting sensitive data in a way that allows arithmetic operations to be performed directly on the ciphertexts. Some privacy issues could arise through reverse-engineering the algorithm or query outcomes [DN10, KM11]. To address this, the differential privacy framework can be used to provide individuals with guaranteed plausible deniability by adding suitably calibrated noise to query outcomes. We will also describe federated learning, which is needed for multiple parties to jointly learn a machine learning model from distributed data. We will start the section with hashing, a foundational technique for many privacy technologies.
3.1 Hashing Techniques
Hashing has a long history in Computer Science. Hashing was first introduced in 1953 to speed up the searching for information in storage systems [Kno75]. The general idea of hashing, or more specifically a hash function, is to map values that are of arbitrary sizes, such as textual strings that can have different lengths, to fixed sized values such as integer numbers that are in a specific range. A hash function therefore converts input values into hash values or hash codes. These hash values are then used as indexes or pointers into a hash table of fixed size. According to [CZ17], a good hash function needs to: (1) map values to hash values as uniformly as possible across the space of possible hash values, (2) the same input value will always result in the same hash value, (3) the mapping process should be fast, and (4) the size of the hash table for a given set of values can be specified by setting the range of integer hash values.
Hashing has been and continues to be a valuable security mechanism for making data unreadable to the human eye, preventing its interception by malicious individuals, and providing a way to validate its integrity. Next we explain few variations of hashing that are useful in the financial intelligence domain.
Keyed Cryptographic Hashing
A keyed hash function (also known as a hash message authentication code, or HMAC) is an algorithm that uses a cryptographic key and a cryptographic hash function to produce a message authentication code that is keyed and hashed [BCK96]. In general HMAC is used to verify the integrity and authenticity of data by computing a fixed-size hash value based on a secret key and input data.
Unlike traditional cryptographic hash functions that generate a hash value solely based on the input data, keyed hashing incorporates a secret key or salt value into the hashing process, making it resistant to tampering and unauthorised modification. The key serves as a shared secret between the sender and receiver, ensuring that only parties possessing the key can compute or verify the hash value. This provides a strong level of security against malicious tampering or forgery attempts, as any alteration to the input data or hash value would require knowledge of the secret key to generate a valid HMAC. Keyed cryptographic hashing is widely used in various security applications, including message authentication, data integrity verification, and secure communication protocols [KSN+23].
Bloom Filters
A Bloom filter [Blo70] is a space-efficient probabilistic data structure used to test whether an element is a member of a set. It is implemented as a bit array $BF[i]$, where $i\in[1,n]$, consisting of $n$ bits. The Bloom filter uses $k$ independent hash functions $H_{1}(x),\cdots,H_{k}(x)$, each of which maps an element $x$ in a set $S$ of $m$ elements to one of the $n$ bits within the bit array $BF$. To determine if an element $x$ is present in the Bloom filter, one needs to check if all $k$ hash functions map $x$ to bits that are set to 1 in the Bloom filter. Because of hash collisions, false positives can occur in a Bloom filter, but false negatives are impossible.
Figure 5: The Dice similarity calculation between the names ‘peter’ and ‘pete’, converted into bigrams ($q=2$) and then encoded into two Bloom filters $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ of length $l=12$ bits using $k=2$ hash functions. The bit “$1$” shown in italics at position $6$ in $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ is a hash collision, because both ‘pe’ and ‘te’ are hashed to this position.
In the privacy-preserving record linkage domain, Bloom filters are commonly used for matching string values or computing similarities between string pairs in a privacy-preserving manner [SBR09, CRS20]. When matching string values, each string is first converted into a set of q-grams, where each q-gram is typically of length 2 or 3. These q-grams are then encoded into a Bloom filter using $k$ hash functions. To compute similarities between Bloom filters, a set-based similarity function such as Jaccard or Dice can be used. Recent literature has also shown that Bloom filters can be used to match both numerical and textual values in databases [WVKR23]. Figure 5 illustrates an example of using a Bloom filter for data matching. The Privacy-Preserving Record Linkage step in Figure 4 can be solved using variations of Bloom Filters. The hash functions in such Bloom Filters are usually keyed cryptographic hashing functions to protect against dictionary attacks.
3.2 Homomorphic Encryption
Homomorphic encryption (HE) is a class of encryption schemes that allows certain arithmetic operations to be performed directly on the ciphertext [AAUC18]. HE schemes are usually public-key cryptosystems. Let $m_{1}$ and $m_{2}$ be two plaintexts, and let $pk$ and $sk$ be the public key and secret key for the encryption Enc() and decryption Dec() functions, respectively. The correctness of a HE scheme is expressed by the following property. For an operation $\diamond$ in the plaintext space, there is a corresponding operation $\bullet$ in the ciphertext space such that
$\text{Dec}(sk,\text{Enc}(pk,m_{1})\bullet\text{Enc}(pk,m_{2}))=m_{1}\diamond m_{2},$ (1)
where the different symbols indicate that the arithmetic operations, either addition or multiplication, in the plaintext and ciphertext spaces may differ. The term “homomorphic” reflects that the decryption function acts like a homomorphism, preserving the structure of the mathematical object.
Depending on the types of arithmetic operations permitted and the number of operations it can handle before decryption fails, an HE scheme is classified into one of the following types, where the algorithmic complexity of the operations we want to perform on data is measured in terms of the depth of circuits and the arithmetic gates supported.
  • Partially HE (PHE) - Schemes that can evaluate circuits containing only one type of arithmetic gate, that is, either addition or multiplication, for unbounded circuit depth.
  • Somewhat HE (SHE) - Schemes that can evaluate a subset of circuits containing both addition and multiplication gates, where the complexity grows as the circuit depth increases.
  • Fully HE (FHE) - Schemes that can evaluate all circuits containing both addition and multiplication gates for unbounded circuit depth.
The privacy-preserving record linkage and model scoring steps in Figure 4 can be done using PHE, but full-scale risk modelling using both supervised and unsupervised machine learning algorithms would usually require SHE or even FHE. FHE was made possible only in recent years due to the breakthrough described in [Gen09]. Gentry’s work is technically based on lattice-based cryptography, an active research area also seen as a strong candidate for quantum-resistant encryption [A+22]. In addition to the standard HE components, Gentry’s scheme introduces two key concepts: squashing and bootstrapping. Simply put, squashing reduces the complexity of decryption by providing a hint to the evaluation process (which introduces additional security assumptions), while bootstrapping allows the scheme to evaluate its own decryption along with an arithmetic operation. When combined, these techniques enable the scheme to perform unbounded arithmetic operations with controlled decryption errors.
It is worth noting that HE schemes are inherently probabilistic, meaning that the same plaintext can produce a different ciphertext each time it is encrypted. This is to ensure a security notion called semantic security. For a more thorough introduction to HE and its mathematical and computer science foundations, the reader is referred to [Hal17, Pei16, LNP22].
To conclude this section, we provide an example of a partially homomorphic encryption (PHE) scheme: the multiplicative HE scheme known as ElGamal Encryption [ElG85]. This scheme is based on the Diffie-Hellman key exchange protocol [DH76] to securely exchange a masked secret key between Alice and Bob for encryption and decryption.
Public key ElGamal Encryption. Bob is the sender, Alice is the receiver. Alice chooses a cyclic group $G$ with order $q$ and a generator $g$.
Private key: Alice’s secret key is a random integer $x\leftarrow\{1,\dots,q-1\}$. Public key: Alice computes $h=g^{x}\in G$ and publishes the public key $(G,q,g,h)$. Encryption: To encrypt a message $m\in G$, Bob randomly selects $y\leftarrow\{1,\dots,p-1\}$, computes the shared secret $s=h^{y}\in G$, and sends the ciphertext $(c_{1},c_{2})$, where $c_{1}=g^{y}$ and $c_{2}=m\cdot s$. Decryption: To decrypt, Alice computes the shared secret $s=c_{1}^{x}$ and its inverse $s^{-1}\in G$, then recovers the message $m=c_{2}\cdot s^{-1}$.
Since each term in the ElGamal ciphertext can be multiplied separately to produce a multiplication ciphertext, and Alice can recover the correct multiplied message using the above decryption process, it implies that ElGamal is a multiplicative HE scheme. Specifically, if $C=(c_{1},c_{2})\leftarrow\text{Enc}(m)$ and $\tilde{C}=(\tilde{c}_{1},\tilde{c}_{2})\leftarrow\text{Enc}(\tilde{m})$, then decryption of the product $C\cdot\tilde{C}$ yields $m\cdot\tilde{m}$, i.e., Dec$(C\cdot\tilde{C})=m\cdot\tilde{m}$. We will see some applications of the additive homomorphic form of ElGamal encryption in Section 4.2.
3.3 Secure Multiparty Computation
Secure multiparty computation (MPC) is a concept that drives the development of protocols designed to securely and accurately perform computations across multiple parties. Research in MPC originated with Yao’s so-called Millionaires’ problem [Yao82], in which two parties seek to know who is richer but without having to disclose their net wealth to each other or a trusted third party. There are many solutions to the Millionaires’ problem involving different cryptographic primitives. Here we briefly describe an accessible solution from [LT05]. The algorithm first converts two numbers in binary representation $x=x_{n}x_{n-1}\ldots x_{1}$ and $y=y_{n}y_{n-1}\ldots y_{1}$ we want to privately compare to the following two sets
$S_{x}=\{x_{n}x_{n-1}\ldots x_{i+1}1\,|\,x_{i}=0,1\leq i\leq n\}$
$S_{y}=\{y_{n}y_{n-1}\ldots y_{i}\,|\,y_{i}=1,1\leq i\leq n\}$
and then securely compute the cardinality $|S_{x}\cap S_{y}|$. The correctness of the procedure comes from the fact that $x>y$ if and only if $S_{x}\cap S_{y}$ is non-empty. There are various private set intersection cardinality (PSI-CA) protocols for computing $|S_{x}\cap S_{y}|$, utilising different techniques such as homomorphic encryption, hashing, Bloom filters, and oblivious transfer. Here, we describe a PSI-CA protocol that employs hashing and commutative encryption from [AES03].
Informally, an encryption function $f$ is commutative if the order of encryption does not matter for arbitrary encryption keys $e$ and $e^{\prime}$, i.e., $f_{e}\circ f_{e^{\prime}}=f_{e^{\prime}}\circ f_{e}$ where $\circ$ is function composition. An example given in [AES03] is the function
$f_{e}(x)=x^{e}\mod p,$
where the powers commute, yielding
$(x^{e}\bmod p)^{e^{\prime}}\bmod p=x^{ee^{\prime}}\bmod p=(x^{e^{\prime}}\bmod p)^{e}\bmod p.$
The security of this commutative encryption is based on the Decisional Diffie-Hellman hypothesis [Bon98]. Armed with a commutative encryption function, $|S_{x}\cap S_{y}|$ can be computed via double encryption as follows:
$|S_{x}\cap S_{y}|=|f_{k_{y}}(f_{k_{x}}(S_{x}))\cap f_{k_{x}}(f_{k_{y}}(S_{y}))|,$
where $k_{x}$ and $k_{y}$ are the encryption keys for the two parties holding the value $x$ and $y$, respectively.
The above secure number comparison protocol provides a glimpse into the rich literature on MPC [Gol98, CD15, EKR+18]. Interestingly, what makes a distributed computation task an MPC protocol cannot be straightforwardly defined in a mathematical way. The well-accepted way of defining MPC’s security is by simulating an ideal world and a real world, known as the ideal/real simulation paradigm [Lin20]. In the simulation, the two worlds are identical except that there is a trusted third party in the ideal world but not in the real world. To examine the security of a protocol under adversarial attacks, one would compare the outcomes of the protocol in the two worlds to determine if the adversary in the real world is able to gather any additional non-disclosed information, should the same attacks also be successful in the ideal world.
With the ideal/real simulation paradigm in place, the following are important properties of an MPC protocol. Privacy guarantees that an adversary learns nothing about an honest party’s input beyond what is prescribed. Correctness ensures that the protocol always produces the correct result despite being attacked by adversaries. Guaranteed output stops the adversary from halting the protocol hence stopping others from receiving an output. Lastly fairness states that an adversary receives the protocol output if and only if an honest party receives the protocol output.
There are two primary adversarial models: semi-honest and malicious adversaries. In the semi-honest model, an adversary correctly follows the protocol specification, but seeks opportunities to learn the protected privacy from information obtained during the computation process. In contrast, the malicious adversarial behaviour can arbitrarily deviate from the protocol, including aborting the protocol or changing its inputs and outputs.
3.4 Differential Privacy
Differential privacy (DP) is a technique that aims to enable the analysis of sensitive data while protecting the privacy of individuals whose data is included in the analysis [DR14]. In general, DP is achieved by adding noise in a controlled manner during the analysis process, ensuring that statistical queries do not reveal information about specific individuals. This allows organisations to derive valuable insights from their data while minimising the risk of unauthorised disclosure or misuse of sensitive personal data.
Definition 3.4.1.
A randomised algorithm $\mathcal{A}$ is $\epsilon$-differentially private if for all neighbouring databases $\mathbf{D}$ and $\mathbf{D^{\prime}}$ that differ in at most one record, and for all sets $\mathcal{O}$ outputs, we have
$Pr[\mathcal{A}(\mathbf{D})\in\mathcal{O}]\leq\exp(\epsilon)\cdot Pr[\mathcal{A}(\mathbf{D^{\prime}})\in\mathcal{O}],$
where $Pr[\cdot]$ denotes the probability of an event.
The parameter $\epsilon$ is known as the privacy budget in DP. A smaller value of $\epsilon$ indicates a stronger privacy guarantee, as it reflects a stricter bound on how much the query results can change in response to the inclusion or exclusion of a single record. The Bayesian semantics of differential privacy can be found in [KS14, KM14, YZN24]; in short, DP guarantees that the odds ratio of an attacker’s prior and posterior beliefs (before and after seeing the query output) in whether a specific record is part of the input database is bounded, and that confers every individual with plausible deniability on whether the person’s data is in the dataset.
An example scenario that illustrates Definition 3.4.1 is differencing attacks. Consider a medical database containing records of individuals who participated in a disease screening. An analyst with access only to counting queries can query the database under different conditions to select two distinct subsets of the data. If the difference between these subsets is designed to isolate a specific individual, the analyst could potentially infer whether or not this individual has the disease by examining the difference between the two query results. Differential privacy prevents this by injecting suitably calibrated noise into the results of these queries, thus preventing the analyst from reliably determining the presence or absence of any particular individual based on the differences in query results.
A fundamental DP technique is the Laplace Mechanism [DMNS06, DR14]. Given a real-valued query function $f:\mathbb{N}^{|\chi|}\rightarrow\mathbb{R}^{k}$, the Laplace Mechanism directly adds Laplace noise to the query outputs as
$M_{L}(\mathbf{D},f,\epsilon)=f(\mathbf{D})+(e_{1},\dots,e_{k}),\text{ where }e_{i}\leftarrow\text{Lap}\left(\frac{\Delta f}{\epsilon}\right),$
where Lap() is the Laplace distribution, and
$\Delta f=\max_{\mathbf{D}\sim\mathbf{D}^{\prime}\in\mathbb{N}^{|\chi|}}{||f(\mathbf{D})-f(\mathbf{D}^{\prime})||_{1}}$
is a measure of the sensitivity of the query on two neighbouring databases $\mathbf{D}$ and $\mathbf{D}^{\prime}$. The independent and identically distributed noise is randomly sampled from the Laplace distribution with mean 0 and scale $\Delta f/\epsilon$. The noise’s scale is proportional to the query sensitivity (when $\epsilon$ is fixed) and inversely proportional to the noise budget $\epsilon$ (when $\Delta f$ is fixed). This makes good sense, since a less sensitive query needs less noise to hide its output variation. A smaller $\epsilon$ indicates stronger privacy guarantee hence more noise needs to be added to the outputs.
In the financial domain, where large volumes of sensitive data are collected and analysed for various purposes such as risk assessment, fraud detection, customer segmentation, and personalised marketing, DP serves as a valuable tool for balancing the need for data-driven insights with privacy protection. By incorporating DP into data analytics workflows, financial institutions can adhere to stringent regulatory requirements, such as GDPR and Basel III [^5] See https://gdpr-info.eu and https://www.bis.org/bcbs/basel3.htm, while still leveraging the richness of their datasets for business intelligence and decision-making [Nis21]. Essentially all financial data are graph data and a survey of DP algorithms for graph data can be found in [LPR+23].
3.5 Federated Learning
Federated learning (FL) [MMR+17] is a new learning paradigm that aims to build a joint machine learning model based on data distributed across multiple sites or owned by different participants. During model training, information such as gradients is exchanged between participants, but the raw data are not shared. The intention is that the exchanged information does not reveal any private data belonging to any party.
Figure 6: A federated learning architecture. In order to guarantee the privacy of the data, federated learning only permits all the data owners to exchange the model gradient ($w$) with the aggregation server. During this process, each data owner trains its own model with the local data, then uploads the local model to the aggregation server. After aggregating all the received models, the server returns the new global model updates ($W$) to each data owner.
Figure 6 illustrates a typical FL process. Initially, each data owner receives a generic global model from the aggregation server. After receiving the initial model, each data owner conducts local training on this model with their own data and then uploads the related gradient information (local model updates) to the aggregation server. The aggregation server averages the updates received from data owners into the global model, then updates the global model to replace each user’s local model. This process repeats until the global model achieves the desired performance or the training reaches the maximum iteration number.
FL can be categorised into three types based on the distribution of data: horizontal FL, vertical FL, and federated transfer learning [LSTS20]. Horizontal FL occurs when the user features of the datasets overlap significantly, but the users themselves overlap minimally. Vertical FL occurs when the user features of the datasets overlap minimally, but the users themselves overlap significantly. In cases where both users and user features of the datasets overlap rarely, transfer learning can be used to address the lack of training data [LKX+20].
Privacy protection is a major concern in FL [ARPS20, NDP+21]. Although FL protects private data on each device by exchanging only model gradients with the aggregation server, the model communication during the entire training process can still potentially leak sensitive information to third parties, for example, through reverse engineering of models [ZLD22]. Two techniques can be employed to prevent private data leakage during FL model transmission:
  • Secure aggregation techniques ensure that raw data or local updates remain private during the aggregation process. Methods such as MPC or HE (described in Sections 3.3 and 3.2) can be used to aggregate updates without revealing individual contributions.
  • DP methods add noise to local updates to protect individual privacy. This noise ensures that the contribution of any single device is indistinguishable from the contributions of others. Aggregating noisy updates provides privacy guarantees while still allowing the central model to learn from the collective data.
In federated learning, model scoring and model learning are two distinct phases of the process. Model scoring refers to the evaluation or assessment of the performance of a machine learning model using a set of evaluation metrics. Common evaluation metrics for model scoring include accuracy, precision, recall, F1-score, and area under the ROC curve (AUC). Both federated model learning and federated model scoring are important steps in the future-state financial intelligence proposed in Figure 4.
4 Privacy-Preserving Information Sharing
In this section, we describe several real-world financial intelligence scenarios wherein information-sharing between different parties is crucial to support effective detection of financial crimes. In each scenario, we discuss the privacy risks during multiparty collaborations and examine the feasibility of applying various privacy-preserving techniques to address these privacy risks.
It is important to recognise that each scenario may entail varying levels of privacy risks. These risks depend on the degree of trust among different parties. The level of trust is basically inversely proportional to the risks of sharing information that have to be shouldered by the collaborating parties, which is in turn a function of the prevailing government legislation (or lack thereof) and applicable penalties in the event of a privacy breach.
The parties involved in conducting financial intelligence analysis today may include both public sectors (domestic or international) and private sectors. Generally speaking, information-sharing between different domestic public sectors involves a high level of trust. In contrast, information sharing between public and private sectors, or among private sectors, often entails a low level of trust. This lack of trust may stem from various factors, including the absence of clear legislation and threats of business competition. For instance, financial institutions tend to refrain from sharing customer transaction data among themselves, despite the potential benefits such information-sharing could offer in enhancing their anti-money laundering controls. Privacy technologies provide the greatest value in low-trust settings that exist either because the parties involved cannot trust each other or a third party for commercial/legal reasons, or when there is ambiguity on whether a public-good sharing of information crosses the threshold into unacceptable privacy intrusion.
4.1 Suspicious Persons List
Problem statement
Suppose an FIU possesses a moderate-sized list of suspicious individuals (or, more generally, commercial entities) obtained, say, from its own investigations and law-enforcement partner agencies and the FIU wishes to acquire additional information about those suspicious individuals, including where they conduct their financial activities and whether they are known to other FIUs. This can be done if the FIU can safely distribute the Suspicious Person List (SPL) to other FIUs and financial institutions to verify the association of individuals on the list with specific financial entities before engaging in further targeted information requests or exchanges with those financial entities.
Privacy risks
If the SPL is shared with the financial institution in a manner that allows for the identification of individuals on the list, there arise the following privacy concerns, especially in the low-trust public-to-private information-sharing setting.
  • (Accidental Tip-off) The financial institution may inadvertently disclose the FIU’s suspicions to an existing customer, prompting the customer to take additional measures to conceal their activities.
  • (Debanking and Displacement) The financial institution may deliberately cease its banking relationship with a suspicious individual before conclusive evidence can be gathered to substantiate or refute the individual’s suspected wrongdoing. This action may be deemed discriminatory if the financial institution’s decision to terminate services is also influenced by personal factors such as the individual’s nationality, race, or name. Further, suspected individuals that are debanked will also move on to other parts of the financial system, including the shadow financial system, and the FIU will lose sight of them.
  • (Need to Know) The FIU can inadvertently disclose the identities of suspicious entities who are not customers of a financial institution, and that would be an unacceptable violation of those individual’s privacy and the need-to-know principle.
Proposed solution
The problem described above of identifying common individuals between the SPL and each participating financial institution’s customer database can be formulated as a private set intersection (PSI) problem. In a two-party computation protocol, the goal of PSI is to identify the common elements that appear in both parties’ sets and nothing else.
We now describe a simple semi-honest PSI protocol from [FNP04] that utilises polynomial evaluation and homomorphic encryption to ensure the protection of unmatched individuals in the SPL. For detailed security and efficiency analyses, readers are encouraged to refer to [FNP04]. For simplicity, it is assumed that each individual can be uniquely identified by an identification (ID) number, which can be derived from a national ID (in some countries) or a combination of personal information. Here are the steps of the protocol.
  1. 1.
    Assuming the FIU’s list of suspicious individuals is $S=\{s_{1},\dots,s_{n}\}$, the FIU first defines the polynomial
    $f(x)=(x-s_{1})(x-s_{2})\cdots(x-s_{n}),$
    whose roots are exactly the elements in the set $S$. The question of whether an arbitrary $s^{\prime}$ is in $S$ can be checked by evaluating $f(s^{\prime})$.
  2. 2.
    The FIU uses polynomial interpolation to find the coefficient representation of $f(x)$
    $f(x)=a_{0}+a_{1}x+\cdots+a_{n}x^{n},$
    which is uniquely defined according to the Fundamental Theorem of Algebra. The FIU then encrypts the coefficients of $f(x)$ using HE and send the encrypted coefficients and the public key to the other parties:
    $\text{Enc}(f)=\{\text{Enc}(a_{0}),\dots,\text{Enc}(a_{n})\}.$
  3. 3.
    Upon receiving the encrypted $f$, each financial institution evaluates the encrypted polynomial $\text{Enc}(f)$ at each element $d\in D$ in its customer database, then multiplies the evaluation result by a random integer $r$ and adds it to $d$. The property of HE guarantees that the result of these operations is equivalent to the following
    $\text{Enc}(rf(d)+d),\forall d\in D.$
    The above formula ensures that if $d$ is a positive match then FIU is able to identify it from the output, and if $d$ is not a positive match then FIU learns nothing about $d$ from the output.
  4. 4.
    Once the matching outputs are shared, FIU decrypts the outputs and find those elements that also appear in its suspicious person list.
By executing this PSI protocol, the FIU and financial institutions collaborate in a confidential manner, enabling the FIU to identify individuals who appear on both the FIU’s SPL and the databases of participating financial institutions. This facilitates the FIU in requesting additional information from appropriate financial institutions to conduct further investigations. Simultaneously, individuals listed on the SPL but absent from a financial institution’s database remain undisclosed to that institution. Similarly, individuals present in a financial institution’s database but not flagged on the SPL remain confidential to the FIU.
Industrial grade solutions
The FIU.net operated jointly by FIUs in the European Union has offered privacy-preserving sharing of suspicious entities for over a decade now, using the Ma3tch algorithm [BM13]. A recent legal review of FIU.net can be found in [Mou20]. In Australia, the Australian Financial Crime Exchange [^6] http://afcx.com.au is a platform for financial institutions to share lists of entities of interest, from a financial crime and cyber crime perspective, with each other and, selectively, with the Australian Taxation Office.
There are now many fast PSI algorithms that can operate at large scale, including [CLR17, RT21, CM20]. A key issue with PSI is that an adversary can use repeated PSI computations to completely recover an organisation’s entire dataset. Proposals like [PLN23] seek to address this issue by introducing DP protection into PSI computations.
4.2 Transaction Flow Tracing and Typology Matching
Problem statement
A key problem in financial crime investigation is tracing the entire digital footprint of “dirty money” within the financial system, allowing investigators to clearly track the movement of funds across different financial institutions, how they are processed and transferred and, crucially, their ultimate usage or destination. We call this problem transaction-flow tracing.
A key task in financial intelligence supported by transaction-flow tracing is the identification of accounts that match certain money-laundering or terrorism-financing (ML/TF) typologies. The NDIS fraud described in [BILC23] serves as an example. Australia’s National Disability Insurance Scheme (NDIS) provides financial support to eligible participants to improve their quality of life. However, continuous fraudulent activities against the NDIS scheme have resulted in several billions of dollars in losses from its intended use [AUS20]. These fraudulent transactions share common patterns, crossing multiple organisations and eventually being transferred overseas. No single institution in the financial system would have all the data required to identify accounts and entities that exhibit such behaviour. We will show in this section how privacy technologies can help tackle this broad class of problems.
Privacy risks
This scenario exemplifies a case of government-to-private and private-to-private information sharing, which are generally low-trust settings. Complex ML/TF typologies can usually be described as a graph pattern, where nodes correspond to accounts/entities and edges correspond to transactions between those nodes and there can be non-trivial conditions attached to the nodes and edges. To enable the identification of entities that match a complex ML/TF typology that span the financial system, we would need accounts and transaction data from multiple financial institutions to be matched at scale before a filtering or sub-graph matching operation is performed. This is obviously highly intrusive to entities who are not involved in financial crimes, so we need a way of performing the large-scale data-matching and graph pattern searching in a privacy-preserving manner.
Proposed solution
The FinTracer solution proposed in [BILC23] is basically a way to perform an efficient and privacy-preserving breadth-first search across a large distributed graph in homomorphically-encrypted space, allowing Financial Intelligence Units like AUSTRAC to work with law-enforcement agencies and financial institutions to start with some known suspicious entities and find their footprint across the financial system
  1. 1.
    without having to disclose the suspicious entities to financial institutions, thus avoiding the risk of accidentally tipping off the entities; and
  2. 2.
    without having to violate the privacy of unrelated entities even though their data need to be processed.
The key observation behind this algorithm is that large distributed graphs can be represented using sparse matrices and graph traversal can be expressed as matrix multiplication operations [Bur21]. This basic graph search algorithm can serve as a building block to detect a large range of financial crime typologies (i.e. graph patterns), thus significantly advancing the state-of-the-art in intelligence agency’s ability to detect complex financial crimes in a responsible, privacy-preserving manner at a large scale.
[…]
Figure 7: A simplified transaction network to demonstrate FinTracer on the NDIS example. Accounts with cross organisation transactions are linked by directed edges. There are four financial institutions A, B, C and D, each having a number of accounts. The account a1 in A directly received NDIS payment.
To see how this works, suppose AUSTRAC is interested in identifying the accounts (from multiple financial institutions) that transferred NDIS funds overseas. These accounts either received funds directly from NDIS or indirectly via a limited number of cross-institution transactions. The FinTracer tag $t_{\leq}$ is implemented to keep track of accounts with at most $k$ cross-institution transactions. These tags are homomorphically encrypted to securely follow NDIS funds in the entire transaction network. The key components of the FinTracer process are the ElGamal encryption scheme [ElG85] over the additive Curve25519 for its performance efficiency [Ber06], and the aggregations of the tag values across financial institutions over multiple iterations of the breadth-first search algorithm.
Figure 7 is a simplified example of the NDIS fraud scenario. In this example, there are four financial institutions. The only account that directly received suspicious NDIS fund is a1 in institution A. All institutions are only aware of the source and destination accounts that are involved in the transactions they process. In its simplest form (Algorithms 1 and 2 [BILC23]), the FinTracer protocol does the following steps:
  • Initialisation: Each financial institution sets the initial FinTracer tag of each of its customer account $x$ to either $\text{Enc}(1)$ or $\text{Enc}(0)$, depending on whether $x$ receive NDIS fund directly. In the example above,
    • $t_{\leq}(a1)=\text{Enc}(1)$ and $t_{\leq}(\{b1,b2,c1,c2,d1\})=\text{Enc}(0)$ (the set notation denotes that the tag is set for element in the set).
    • The other accounts do not have cross-institution transactions, hence are out of considerations by the protocol.
  • Iteration 1: In the first iteration, the initialised tags are passed to the neighbouring institutions following the transaction flow. The receiving institutions then aggregate the tags with the received tags. In the example, the shared partial mappings are
    • A $\rightarrow$ B: $P^{A}_{B}((a1,b1))=\text{Enc}(1),P^{A}_{B}((a1,b2))=\text{Enc}(1)$.
    • B $\rightarrow$ C: $P^{B}_{C}((b1,c1))=\text{Enc}(0),P^{B}_{C}((b2,c2))=\text{Enc}(0)$.
    • C $\rightarrow$ D: $P^{C}_{D}((c1,d1))=\text{Enc}(0),P^{C}_{D}((c2,d1))=\text{Enc}(0)$.
    After receiving these partial mappings, the aggregated tags are
    • $t_{\leq}(\{a1,b1,b2\})=\text{Enc}(1)$ and $t_{\leq}(\{c1,c2,d1\})=\text{Enc}(0)$
    since $t_{\leq}(b1)=t_{\leq}(b1)+P^{A}_{B}((a1,b1))=\text{Enc}(0)+\text{Enc}(1)=\text{Enc}(1)$ and similarly for $t_{\leq}(b2)$. The tag values of $c1$, $c2$, $d1$ remain $\text{Enc}(0)$ since the initial and received tags are $\text{Enc}(0)$.
  • Iteration 2: The partial mappings are updated by institution B yielding
    • A $\rightarrow$ B: $P^{A}_{B}((a1,b1))=\text{Enc}(1),P^{A}_{B}((a1,b2))=\text{Enc}(1)$.
    • B $\rightarrow$ C: $P^{B}_{C}((b1,c1))=\text{Enc}(1),P^{B}_{C}((b2,c2))=\text{Enc}(1)$.
    • C $\rightarrow$ D: $P^{C}_{D}((c1,d1))=\text{Enc}(0),P^{C}_{D}((c2,d1))=\text{Enc}(0)$.
    After receiving these partial mappings, the aggregated tags are
    • $t_{\leq}(a1)=\text{Enc}(1)$, $t_{\leq}(\{b1,b2\})=\text{Enc}(2)$, $t_{\leq}(\{c1,c2\})=\text{Enc}(1)$ and $t_{\leq}(d1)=\text{Enc}(0)$.
  • Iteration 3: The partial mappings are updated by B and C yielding
    • A $\rightarrow$ B: $P^{A}_{B}((a1,b1))=\text{Enc}(1),P^{A}_{B}((a1,b2))=\text{Enc}(1)$.
    • B $\rightarrow$ C: $P^{B}_{C}((b1,c1))=\text{Enc}(2),P^{B}_{C}((b2,c2))=\text{Enc}(2)$.
    • C $\rightarrow$ D: $P^{C}_{D}((c1,d1))=\text{Enc}(1),P^{C}_{D}((c2,d1))=\text{Enc}(1)$.
    After receiving these partial mappings, the aggregated tags are
    • $t_{\leq}(a1)=\text{Enc}(1)$, $t_{\leq}(\{b1,b2\})=\text{Enc}(3)$, $t_{\leq}(\{c1,c2\})=\text{Enc}(3)$ and $t_{\leq}(d1)=\text{Enc}(2)$.
The FinTracer tags are now sent to the FIU and decrypted. Only those accounts with a non-zero tag have received money that can be traced either directly or indirectly to suspicious NDIS payment.
Industrial grade solutions
An open-source, scalable implementation of FinTracer can be found at http://github.com/AUSTRAC. The implementation includes a C/Cuda library for additive HE with ElGamal over the Ed25519 elliptic curve. A number of algorithms for different graph operations are included in http://github.com/AUSTRAC/ftillite/Documentation/Purgles. The extension of FinTracer, which allows only linear operations because of the use of additive HE, to non-linear set and graph operations to allow the detection of more-or-less arbitrarily complex criminal typologies is described in [BCF23]. Note that differential privacy noise is added in several steps in these algorithms to avoid variations of database-reconstruction attacks [DN03, DMT07].
Finding key players in transaction networks
Once the small number of accounts and entities that matched a certain ML/TF typology is detected using a technology like FinTracer, an FIU like AUSTRAC has the legal power to request much more comprehensive information about those accounts and entities from financial institutions because the request for information or notice to produce information is now targeted and backed by genuine grounds for suspicion. (The remaining small-scale privacy concerns can mostly be addressed using differentially private graph search algorithms like [KRWY16].) With such a dataset, it then makes sense to calculate connected components and risk metrics like PageRank scores and betweenness centrality scores, some of which have been shown to be useful features to enhance the accuracy or confidence of machine learning or rule-based financial intelligence classification models [MCF+17, SvHA+19, vEDvdB+24].
4.3 Collaborative Model Learning
Problem statement
The use of supervised learning and anomaly detection algorithms to learn financial crime models from data has long been a topic of interest within financial intelligence [SvHA+19, vEDvdB+24], but the practical utility of such approaches has been hampered by a lack of quality labelled data in the case of supervised learning algorithms, and high false positive rates in the case of anomaly detection algorithms.
Recent advances in privacy-preserving FL have the potential to alleviate some of these issues by allowing different business units within a financial institution and even multiple financial institutions to jointly learn models from their collective data. In this section, we describe some of the key ideas behind some exploratory work in this area.
Privacy risks
To enable FL, we would need accounts and transaction data from multiple financial institutions to be matched, often probabilistically, at scale before a machine learning algorithm can be run. In the case of supervised learning, the participating parties would also need to share, as labelled data, entities that are either known or suspected to be engaging in financial crime. This is obviously highly intrusive to entities who are not involved in financial crimes, and also causes the kind of tipping-off issues discussed in Section 4.1 for those (suspected to be) involved in financial crimes. All these imply that we need scalable and secure ways to perform large-scale probabilistic data matching and model learning in a privacy-preserving manner. The publication and reuse of machine learning models by multiple organisations also open the door to attacks like membership inference that need to be managed [MPP+21].
Proposed solutions
As discussed, the first step in FL is privacy-preserving data matching. Unlike transaction flow tracing, where the source and target of each transaction are uniquely identifiable for exact matching, there can be inconsistencies in how different financial institutions collect and store individual information, such as names, addresses, driver’s licences, business names, and other details. Consequently, data matching and entity resolution in federated learning – both within and across financial institutions – are often performed probabilistically. A comprehensive description of key privacy-preserving data matching/record linkage algorithms, including those that are built on Bloom Filters as described in Section 3.1, can be found in [CRS20].
We now describe a simplified example of federated learning to illustrate its applicability in financial intelligence. The task is to learn a linear regression risk model using probabilistically matched data from multiple participating parties, while preserving privacy in a distributed manner. The target of the regression is the individual risk score, and the features for the regression are derived from all the data held by the REs; the setup is shown in Figure 8 for three parties.
More formally, given a training set $\{({\bf\it x}_{l},y_{l})\}_{l=1,\dots,m}$ that includes $m$ individuals, where the feature set ${\bf\it x}_{l}={\bf\it x}_{l}^{(1)}\cup\cdots\cup{\bf\it x}_{l}^{(p)}$ is the union of the features from all parties, the task is to fit a linear regression model
${\bf\it Y}={\bf\it X}{\bf\it w}+{\bf\it\epsilon}$
that predicts the target variable ${\bf\it Y}$ as accurate as possible.
Figure 8: Federated learning setup for privacy-preserving risk modelling
The regression training is performed using gradient descent. The weight vector ${\bf\it w}$ is initially assigned random values and then updated using the gradient of the loss function with respect to each weight element $w_{j}$ as follows, until convergence.
$w_{j}:=w_{j}-\alpha\frac{1}{m}\sum_{l=1}^{m}(\langle{\bf\it x}_{l},{\bf\it w}\rangle-y_{l})x_{l,j}$ (2)
The algorithm converges to the optimal solution because the least-squares loss function is convex and the learning rate $\alpha$ controls the convergence rate.
Below are the steps for the privacy-preserving distributed version of the algorithm that uses an additive HE system.
  1. 1.
    The FIU generates a public-private key pair ($pk$, $sk$) and shares the public key $pk$ with the REs.
  2. 2.
    Each RE initialises its portion of the weight coefficients ${\bf\it w}^{(i)}$ randomly and computes a partial prediction for each individual
    ${\bf\it Y}^{(1)}$ $=(\langle{\bf\it x}^{(1)}_{1},{\bf\it w}^{(1)}\rangle,\langle{\bf\it x}^{(1)}_{2},{\bf\it w}^{(1)}\rangle,\ldots,\langle{\bf\it x}^{(1)}_{m},{\bf\it w}^{(1)}\rangle),$
    ${\bf\it Y}^{(2)}$ $=(\langle{\bf\it x}^{(2)}_{1},{\bf\it w}^{(2)}\rangle,\langle{\bf\it x}^{(2)}_{2},{\bf\it w}^{(2)}\rangle,\ldots,\langle{\bf\it x}^{(2)}_{m},{\bf\it w}^{(2)}\rangle),$
    $\vdots$
    ${\bf\it Y}^{(p)}$ $=(\langle{\bf\it x}^{(p)}_{1},{\bf\it w}^{(p)}\rangle,\langle{\bf\it x}^{(p)}_{2},{\bf\it w}^{(p)}\rangle,\ldots,\langle{\bf\it x}^{(p)}_{m},{\bf\it w}^{(p)}\rangle),$
    where ${\bf\it x}^{(i)}_{l}$ is the feature set of individual $l$ held by the $i^{th}$ RE.
  3. 3.
    Each RE homomorphically encrypts its partial prediction element-wise and participates in a round robin communication to aggregate a complete prediction for each individual. The resulting encrypted prediction is then sent to the FIU for subsequent computation.
    $\overline{{\bf\it Y}}=\left(\sum_{i=1}^{p}\text{Enc}(pk,{\bf\it Y}^{(i)}_{1}),\dots,\sum_{i=1}^{p}\text{Enc}(pk,{\bf\it Y}^{(i)}_{m})\right),$
  4. 4.
    The FIU decrypts the complete prediction and computes the difference from the actual values
    ${\bf\it E}=\text{Dec}(sk,\overline{{\bf\it Y}})-{\bf\it Y},$
    and then sends the unencrypted ${\bf\it E}$ back to the REs.
  5. 5.
    Each RE then updates its portion of the weight vector ${\bf\it w}^{(i)}$ independently using Equation 2
    ${\bf\it w}^{(i)}_{j}={\bf\it w}^{(i)}_{j}-\alpha\frac{1}{m}\sum_{l=1}^{m}{\bf\it E}\cdot x^{(i)}_{l,j}$
    where $j$ ranges over the variables held by the RE.
  6. 6.
    Repeat steps 2 to 5 until the model has converged.
Throughout the entire computation, the FIU only sees the predicted values from the REs, from which it cannot recover the data held by the REs. Conversely, the REs only see the encrypted partial predictions from each other and the gradients from the FIU, from which they cannot recover the data held by the FIU or other REs.
In the remainder of this subsection, we provide a survey of some innovative use of FL that holds promise for financial intelligence. In the work of [SZK+22], we see a reasonably compelling demonstration of how FL can significantly enhance the detection capabilities of financial crime models. Through a series of case studies and rigorous experimental evaluations, this research highlights the dual achievement of improved detection performance and robust data privacy. The study underscores how FL can elevate the precision of financial crime detection while safeguarding sensitive information.
Building upon this foundation, [DNM+23] introduces a novel approach that advances secure and collaborative model training across multiple financial institutions. By integrating HE with FL, this approach achieves secure model aggregation and parameter updates. The experimental results and evaluations presented reveal the approach’s effectiveness in preserving data privacy and security, thus fostering a collaborative environment for financial data analysis across institutions.
Further extending this narrative, [HHIL+17] presents an end-to-end scheme designed for linear classifiers, employing additive HE to counteract threats from semi-honest adversaries in vertical FL scenarios. This research delves into the impact of entity resolution on learning performance, utilising a Taylor approximation for the loss and gradient functions. The application of Paillier additive HE within this context enables privacy-preserving computations, showcasing an innovative method for handling sensitive data securely.
Additionally, [BP20] proposes a FL scheme tailored for fraud detection. In this scheme, individual financial organisations enhance their model parameters with DP noise before transmitting them to a central server. This strategy effectively diminishes the server’s ability to recover private information, advancing the cause of privacy in fraud detection efforts.
Though these FL techniques have yet to achieve industrial-grade maturity, they collectively offer a glimpse into the forefront of financial intelligence. Their contributions highlight ongoing advancements and set the stage for future innovations in the field.
Industrial grade solutions
There are several industrial-grade solutions like [ZCN17, ZNCC18] for conducting probabilistic matching of databases with millions to billions of entities and these can be used in conjunction with techniques like Bloom filters and HE to produce somewhat scalable privacy-preserving probabilistic data-matching algorithms. While federated learning is now a mature technology, progress in using federated learning at the industrial scale for financial intelligence applications has been slow because, in most cases, there is not an abundance of labelled data to support the acquisition of robust and trusted machine-learning models with low false-positive rates. With further advances in FL and a general uplift in the maturity of FIUs in adopting privacy technologies, we expect to see more progress in this area in the future.
5 Discussion and Conclusion
Financial intelligence equips law enforcement and financial institutions with sophisticated analytical tools and data-driven methodologies to proactively detect and prevent financial crimes. While financial intelligence plays a crucial role in supporting Anti-Money Laundering (AML) and Counter-Terrorist Financing (CTF) regimes by delivering actionable intelligence and insights, it may present challenges in relation to data privacy regulations such as Australia’s Privacy Act and European Union’s General Data Protection Regulation (GDPR). These challenges arise particularly when there is a discrepancy between the initial purpose for which data was collected by financial institutions and its subsequent use in various financial intelligence applications. While public and private sectors, industry associations and regulatory bodies are actively working to harmonise AML/CTF and data privacy regulations, such changes often involve lengthy and uncertain political approval and implementation processes.
From the practical perspective of being able to do something about complex financial crimes today, we advocate for the integration of privacy-preserving technologies within financial intelligence tools, not just as a mitigation strategy while waiting for legal experts to navigate the complex AML/CTF reform landscape [Bug18], but to inform those AML/CTF reform discussions. This approach offers several benefits. First, privacy-preserving technologies can be swiftly implemented, allowing financial institutions and financial intelligence units to address data privacy concerns while fulfilling AML/CTF obligations. Second, adopting privacy-preserving methods helps institutions build trust with customers and stakeholders by demonstrating a commitment to both data protection and effective financial crime prevention. Third, privacy-preserving solutions offer a flexible framework that can adapt to evolving regulatory requirements, helping institutions stay compliant as laws and regulations are updated.
We hope this paper enhances public understanding of privacy-preserving financial intelligence through the various AML/CTF use cases we explore. The use cases highlight the evolving nature of financial crimes and the necessity for ongoing innovation in privacy technologies to address emerging threats effectively. Having said that, for financial institutions and regulatory bodies, adopting privacy technologies involves careful consideration of both technological and operational aspects. Effective implementation strategies must balance the trade offs between data privacy and the quality and efficiency of analysis. While these technologies address immediate challenges, they are not a substitute for comprehensive regulatory reforms. Continuous refinement of regulations is essential to align financial crime prevention efforts with data privacy goals.
Acknowledgements
We would like to thank the many former colleagues in AUSTRAC and financial crime units in major Australian financial institutions who spent many hours teaching us the intricacies of the financial intelligence business. We are also grateful to many of them for continuing to play an important role to keep our financial system and our country safe.